OpenStudy (anonymous):

right a quadratic equation in the variable x having the given numbers as solutions. type the equation in standard form, ax^2 + bx + c = 0 solution 7, only soultion.

OpenStudy (asnaseer):

if a quadratic has just one solution x=a, then it can be written in the form:\[(x-a)^2=0\]use this to solve.

OpenStudy (asnaseer):

in general if it has two solution, x=a and x=b, then it can be written as:\[(x-a)(x-b)=0\]

OpenStudy (asnaseer):

so you just need to expand the expression to get it into the form asked for by the question.

OpenStudy (anonymous):

what does it mean when it ssays solution 7 is the only soloution

OpenStudy (asnaseer):

it means x=7 is the only solution

OpenStudy (asnaseer):

so it must be a solution of the quadratic:\[(x-7)^2=0\]as you can see, only x=7 will give this a value of zero. so just expand this expression to get your answer.

OpenStudy (asnaseer):

as another example, suppose they asked for a quadratic that had solutions of 2 and 3. then you would use:\[(x-2)(x-3)=0\]